Felix will receive 20,000 British pounds one month from now as payment for apple juice produced and

export by his company. Felix is concerned about his exposure because he

believes that there are two possible scenarios: (1) the pound will depreciate by 3%

over the next month or (2) the pound will appreciate by 2% over the next month.

There is a 70% chance that Scenario 1 will occur. There is a 30% chance

that Scenario 2 will occur.

Felix notices the spot rate of the pound is GHS 8.1 and the one month

forward rate is GHS 8.6. Felix can purchase a put option over the counter

from a securities firm that has an strike price of GHS 8.6, a premium of

GHS0.025, and an expiration date of one month from now. Determine the amount of

cedis received by the apple Company under each of the two exchange rate

scenarios if:

a) The receivables to be received in one month are not hedged.

b) A put option is used to hedge the receivables in one month.

c) A forward hedge is used to hedge the receivables in one month.
1
Expert's answer
2022-02-21T16:40:03-0500

a)(1) the pound will depreciate by 3%:


"20 000-0.03\times20000=19400"


(2) the pound will appreciate by 2% over the next month:

"20000+0.02\times20000=20400"


b)in the event of a depreciation, the investor will be obliged to fulfill obligations under the contract, under which he will buy from the counterparty at the price at which he planned in advance.

the spot rate of the pound is GHS 8.1

Felix will receive 20,000


In the event of an increase in the market value, the option will not be presented for exercise, and additional profit will appear in the form of an option premium.

8.1+0.025=8.125

Felix will receive 20,000+option premium.


c)forward rate is GHS 8.6

1) the pound will depreciate by 3%:

"8.1-0.03\times8.1=7.857"

8.6-8.1=0.5

guaranteed to receive a profit that can no longer rise or fall.

20 000+0.05

(2) the pound will appreciate by 2% over the next month:

"8.1+0.02\times8.1=8.262"

8.6-8.1=0.5

guaranteed to receive a profit that can no longer rise or fall.

20 000+0.05